OpenStudy (anonymous):

the region by the line y=x/2. teh x axis, and x=2, is rotated around x axis. find the volume of rotation using disk and shell method.. help

OpenStudy (anonymous):

please dont be mean

OpenStudy (inkyvoyd):

Ok, fine. Try wolfram alpha.

OpenStudy (anonymous):

i dont know how to ask the question on that website

OpenStudy (inkyvoyd):

My bad, wolfram won't do it. http://www.khanacademy.org/math/calculus/v/solid-of-revolution---part-2 seriously gives you the answer to sqrt x around x axis.

OpenStudy (anonymous):

buts its x/2

OpenStudy (inkyvoyd):

Ok, my bad again.

OpenStudy (inkyvoyd):

Can you give me a moment? I'll try to do it. I'm 15 years old though, so xDD

OpenStudy (inkyvoyd):

\[\int\limits_{a}^{b} \pi f(x)^2 dx\] is what you are looking for

OpenStudy (inkyvoyd):

i.e, \[\int\limits_{0}^{2}\pi x^2/4\]

OpenStudy (anonymous):

the second is what i got

OpenStudy (anonymous):

what about using the shell metohd

OpenStudy (inkyvoyd):

now, you should get\[(\pi)x^3/12|_{0}^{2}\]

OpenStudy (inkyvoyd):

I don't have any idea. I'm learning calc rite now. Try, r/cheatatmathhomework (reddit.com)